OpenStudy (anonymous):

Solve: 49x^2 – 28x = -4

OpenStudy (anonymous):

That's not quite right Nathan Lets see, you can simplify it by: to (7x-2)(7x-2)=0 so from there solve for x 7x-2=0 7x=2 x=2/7